# Multi-Modal Diffusion for Hand-Object Grasp Generation

### v1

@@ 4.2 Object-Conditioned Grasp Generation
We select the objects from OakInk-Shape test sets to generate grasps on them. Our method can generate visually plausible grasp configurations even though the model has not seen these objects during training.
We perform the quantitative evaluation on the OakInk-Shape test set with the previously introduced metrics. We follow the previous practice [33] to use the widely adopted GrabNet [26] and its refined version GrabNet-refine [33] as the baseline models. We train the models on the GRAB and OakInk-shape train splits. The results are shown in Table 1. Compared to the baseline methods, our proposed method achieves better generation quality per four metrics: FID, Mesh Penetration Depth, Intersection Volume, and Mean Simulation Displacement. Significantly, it achieves nearly 50% improvement regarding the penetration depth. Per the plausibility and truthfulness rated by invited participants, our method also achieves a better rating. Our method achieves a convincingly good performance. Moreover, when we use the objects generated by our method itself, its performance advantage over GrabNet is more significant. This is probably because GrabNet is harder to generalize to object shapes that are significantly different from its training data. On the other hand, as our method learns more generalizable and diverse object prior information by modeling objects in a pure geometric perspective, its robustness to diverse object shapes is much better.
@@ 4.3 Unconditional Grasp Generation
Unconditional grasp generation is a main contribution of this work. Without the massive scale of full-stack datasets such as synthetic robotic manipulation datasets [30], our method is still capable of doing this. It learns the joint distribution between the latent representation of hand and object and leverages the large-scale object datasets to learn a sufficiently generalizable object shape representation model. Compared to the objects generated by the original LION, our method generates objects that are more proper for hand-object interaction after fine-tuning the object prior distribution by the object samples from the selected datasets. We selected some samples from the unconditional grasp generation in the supplementary. We also conduct the quantitative evaluation for the unconditional generation quality by MGD. The results are shown in Table 1. The baseline GrabNet suffers from generating plausible grasps on the out-of-distribution object shapes while our method achieves good results. Moreover, it even achieves on-par performance compared to the generation quality of the baseline GrabNet in the object-conditioned setting, which is very impressive. By learning from large-scale object datasets, the ability of our method to generate object shapes and hand grasp at the same time gives a unique advantage. We think this is a key capacity for the universal grasp generation. Compared to the object-conditioned generated grasps from MGD, the quality of unconditional generation is inferior, which is in fact as expected. This reveals a limitation of existing hand-grasp synthesis methods in many situations. Because the HOI datasets, e.g., OakInk and DexGraspNet [30], usually contain objects from certain categories, with similar scale, geometry, and affordance, the method trained on the training split can learn significantly biased in-domain knowledge about the objects in the test splits. However, a typical object generation model, which is trained on much larger-scale and more diverse object shape datasets, can generate objects with significantly different scales, geometries, and topologies. To generate hand grasps on these out-of-domain object shapes would be a challenge to all existing methods. By designing the method capable of training with object-only data and disentangling the modality representations in the latent diffusion, our method not only has the advantage of generating the object and the hand grasp simultaneously but also achieves much better robustness to generate grasps on out-of-domain objects.
@@ 4.4 Ablation Study
During training, we fuse multiple datasets with heterogeneous annotations to gain as rich object shape and grasp data as possible. This shows the flexibility of our method to be trained with different supervision. However, it also introduces additional variance and potential noise to measure the performance contribution from our methodology. Therefore, we provide ablation studies here to better understand the method’s performance.
@@
Ablation of grasp generation on unseen objects. To provide transparent experimental conclusions, we ablate the training data on the model generation quality. Because OakInk-train data is used for training, which can be biasedly similar to the OakInk-test split, we use the objects from ARCTIC [7] to measure the conditional generation quality on unseen objects. The quantitative evaluation results are shown in Table 2. Compared to using only OakInk-Shape data for training, adding GRAB data and object shapes from Object-only datasets both improves the generalization qualities. The improvement from extra GRAB data is easy to explain as training with more data usually improves the generalizability of the model. It is interesting that adding the object-only datasets also boosts the performance. We believe this is because the extra object data improves the robustness and generalizability of the object encoding and representation modules. It improves the latent representation that the diffusion model learns to generate.
@@
Ablation of unconditional generation quality. The ability to generate both hand and object to form a grasp is another main contribution. We are interested in whether it can benefit from the additional training data as well. We present the ablation study in Table 3. Similar to the conditional generation ablation study results, the unconditional generation quality also benefits from additional training data consistently. The metrics of Pene. Dep., Intsec. Vol. and Sim. Disp. Mean indicate that the generated hand-object pair improves plausibility along with adding more training data. The increasing user scoring result suggests that the generated object and hand also become more and more visually realistic along with adding more training data.
@@
Ablation of synchronous denoising schedulers. As one of the main implementation innovations of MGD, we conduct an ablation about Asynchronous Denoising Schedulers in Table 4. Intuitively, we designed this module to decouple the noise level in hand latent and object latent so that the model could better use the data with heterogeneous annotations and learn to denoise a certain modality. Without Asynchronous Denoising Schedulers, the noise diffusion and denoising time step of the object part is always correlated to the hand latent code. It makes the training biased to the HOI dataset where both modalities are available for supervision. Adding Asynchronous Denoising Schedulers, the denoising is learned with more independence for the object part and the hand part. The model can better learn object shape coding to generate hand grasps. Therefore, when evaluating on the out-of-domain objects generated by MGD, the generation quality is improved as expected.
@@ 5 Conclusion
In this work, we propose a multi-modal grasp diffusion model. By encoding the modalities of objects and hands into a unified latent space, we could generate them either jointly or conditionally. We reveal the limitations of training and test generation quality within the domain and explore how to learn grasp generation on universal objects. Our model can benefit from data with partial annotation thus relieving the limitation of full-stack annotated 3D hand-object grasp data. Neither auxiliary information such as category labels or text descriptions nor optimization is needed in our proposed method. The generated grasp in both unconditional and conditional settings shows good quality.
Appendix A Qualitative Results
We showcase the qualitative visualization results of our proposed method here. We select the object shapes from GRAB and OakInk-Shape that are not used in training as the object condition. The generated hand grasps are presented in Figure 3 and Figure 4. On the same object or objects with similar shapes, we can observe different hand grasp patterns, which are visually plausible. Given that these object shapes are unseen during training and no text/category description is provided as an extra condition, we believe such grasping generation is made from the model’s ability to generate grasping from a pure geometric perspective of the object shape. Moreover, by sampling noise from the latent space, we could generate the object shape and a corresponding hand grasp unconditionally. The results are presented in Figure 5. We could notice the good plausibility of the generated hand-object grasp. The hand pose and articulation are well adapted to the object shapes generated. Compared to the object shapes from real-world scanning, the objects generated by our method lack the same level of well carving. This is also related to the surface reconstruction algorithm, i.e., Alpha Shape [6], which we leveraged to reconstruct mesh from point clouds. Applying a more advanced surface reconstruction method can improve the details and fidelity of the object shapes but that is out of the scope of this work. Moreover, as we study the object shapes from a pure geometric perspective, we can’t use the surface reconstruction arts that require category prior.
Figure 3: Samples of object-conditioned grasp generation on GRAB.
Figure 4: Samples of object-conditioned grasp generation on OakInk-Shape.
@@ Appendix A Qualitative Results
Appendix B Grasp Diversity
Given a single object shape, we could generate a set of different grasps onto it. Though diversity is expected to be constrained by the hand grasp data available in training, there are some other implicit but more fundamental cues to help generate diverse hand grasp. This underlying but fundamental constraint can be leveraged to enhance the generation robustness over unseen objects if the feature of the object shape is sufficiently generalizable. For example, the model can learn to avoid penetration between the object volume and the hand shape and certain fingers should be close to the surface of objects to form a physically valid grasp. We generate grasps with unseen objects from the OakInk-shape test set as the condition in Figure 6. We observe some grasp patterns not provided in the training set, for example, holding a dagger between two fingers. As the implicit concept of forming a valid grasp is always conditioned to object shapes, our strategy of exposing the model training to more diverse object shapes can help to learn grasp patterns on unseen and even out-of-domain objects. We believe this is a key reason for making our method outstanding when generating grasps on out-of-domain objects.
@@ Appendix B Grasp Diversity
Appendix C Dataset Statistics
Here we provide more details about the statistics of the datasets involved in our training and evaluation in Table 5. The HOI datasets with deformable and articulable hand models, i.e., MANO [23], face severe limitations of object resources. Combining the training set of GRAB [26] and OakInk-Image [33] datasets still make just $\sim$100 objects. On the other hand, AffordPose [12] and DexGraspNet [30] contain more object shapes but the different choices of hand models make them hard to integrate with MANO-parameterized datasets. Fortunately, we have the large-scale 3D object shape dataset ShapeNet [4] with more than 50,000 objects. We combine the objects from these datasets in the training for the object part. They help to construct a more universal prior distribution for object generation and the grasp posterior distribution.
Appendix D Grasp-conditioned Image Generation
We demonstrate using MGD as a creativity tool to generate 3D grasps and the corresponding images. In Figure 7, we first use MGD to sample interaction pairs of 3D hands and objects. With the rendering of the 3D hand-object grasp as the input, we applied Adobe Firefly [2], an existing tool to generate the images with the depth and edge reference of the 3D pairs. The hand pose and the geometry provide guidance for Firefly to complement the details given simple text prompts. In Figure 8, compared to grasping images authorized purely by text prompts, images guided by MGD’s output have more consistent gestures, better-aligned object geometry and boundaries, and fewer hand artifacts. Therefore, with the rise of text-to-image generation, our hand-object grasp generation model can serve as a proxy to enhance the consistency among multiple instances and the visual plausibilities. We also provide a closer look at their detailed in Figure 9 to compare the results with and without the generated grasp as a proxy. At each row, the images are generated from the same text prompt as shown at the bottom. There are two main benefits of generating images conditioned to the grasp. It first allows a set of images with aligned hand and object geometry and boundaries which can be useful for image and video editing. On the other hand, even though the Adobe Firefly generator has shown a significant advantage over the public Stable Diffusion in eliminating artifacts, we still observed many finger artifacts when generating the images without a grasp condition. Fortunately, with the generated grasp, including the object shape and the hand, as the condition, the image generator can produce significantly fewer artifacts, especially artifact fingers, which have been a notorious issue in image generation recently. We provided some zoomed-in examples at the bottom. Moreover, compared to previous works using hand shapes to guide image generation [35, 17], we could generate both the object and the hand. Therefore, we could use the whole scene of hand-object interaction as the condition and thus enable more controllable details in the generated image.
Appendix E Failure cases
@@
To provide a transparent evaluation of our proposed method, we still notice some failure cases on the unseen objects from OakInk and GRAB datasets as shown in Figure 10. There are in general three patterns of failures: (1) the hand and the object have no contact, making the grasp physically implausible; (2) the hand skin is twisted or the pose is not able to grasp the object in human common sense; (3) the mesh of object and hand has penetration and intersection. There can be some reasons that make these happen. First of all, we do not explicitly supervise the contact between the object surface and hand as this is a rare annotation on many data sources, and calculating it can be computationally expensive. On the other hand, some implausible grasps have pretty good visual quality as shown in the middle two columns in the figure. However, according to our life experience, we know that the grasp is not physically plausible to manipulate the object. This has gone beyond our scope in this work as we do not have any physics-aware supervision such as the physical demonstration in a physics simulator. Finally, we use the object point cloud to represent object shapes to allow grasp generation on universal objects without a template, this causes the potential penetration between object mesh and hand mesh as the object mesh is ambiguous and unknown to our method. Despite the failure cases, we note that they make only a very small portion of the generated samples (less than 10% by a rough estimation). We show them here for transparency and to help discussion about future works in this area.

### v2

@@ 4.2 Object-Conditioned Grasp Generation
Our method can generate visually plausible grasp configurations even though the model has not seen these objects during training. We generate samples on the objects from GRAB and OakInk test sets as shown in Figure˜3 and Figure˜4 . We perform the quantitative evaluation on the objects unseen during training with the previously introduced metrics. We follow the previous practice (Yang et al., 2022) to use the widely adopted GrabNet (Taheri et al., 2020) and its refined version GrabNet-refine (Yang et al., 2022) as the baseline models. We train the models on the GRAB and OakInk-shape train splits. The results are shown in Table˜1.
Compared to the baseline methods, our proposed method achieves better generation quality per four metrics: FID, Mesh Penetration Depth, Intersection Volume, and Mean Simulation Displacement. The performance advantages are demonstrated on objects from the OakInk test set, generated by our method or ARCTIC dataset.
OakInk training set and training set contain similar objects though not the same. Such a biased similarity between the training and test sets exists in many HOI datasets and conceals the limitation of generating grasps on unseen objects. So we also test on the ARCTIC dataset, which is more confidently out-of-distribution from training. GrabNet-Refine fails to generate grasps with decent quality while the advantage of our method becomes more significant. GrabNet-Refine’s generation quality is inferior to our method with a much larger margin by all metrics. The experiment reveals that some existing methods, such as GrabNet, face difficulty in generalizing to object shapes that are significantly different from training data. On the other hand, our method learns more generalizable and diverse object prior information and more universal grasp generation ability.
Figure 3: Generated samples on objects from GRAB (Taheri et al., 2020) dataset.
Figure 4: Generated samples on objects from OakInk (Yang et al., 2022) dataset.
@@ 4.3 Unconditional Grasp Generation
Our proposed method learns the joint distribution of the latent representation of hand and object. It leverages large-scale object datasets to learn a sufficiently generalizable object shape embedding. To measure the quality of unconditional hand-object grasp, the results are shown in the last row in Table˜1. The metric scores indicate that the unconditional hand grasp generation quality is also good and close to the generation quality on ARCTIC objects. Compared to the object-conditioned generated grasps from JHOD, the quality of unconditional generation is inferior, which is in fact as expected. Here, we also provide the evaluation results by GrabNet-Refine on the same set of objects generated by our method for reference. And GrabNet-Refine fails to generalize to this set of unseen objects again just as its failure on ARCTIC objects. Such failure of generalizability is previously under-explored because the training/test splits of a HOI dataset usually contain objects from certain categories, with similar scale, geometry, and affordance. The method trained on the training split can learn significantly biased in-domain knowledge about the objects in the test splits. To generate hand grasps on more out-of-domain object shapes would be a challenge to all existing methods.
By designing the method capable of training with object-only data and disentangling the modality representations in the latent diffusion, our method not only has the advantage of generating the object and the hand grasp simultaneously but also achieves much better robustness to generate grasps on out-of-domain objects.
@@ 4.4 Ablation Study
We now provide ablation studies to show the contributions of different resources to our method’s final performance.
@@
Ablation of grasp generation on unseen objects. To provide transparent experimental conclusions, we ablate the training data on the model generation quality. We use the objects from ARCTIC (Fan et al., 2023) to measure the conditional generation quality on unseen objects. The quantitative evaluation results are shown in Table˜2. Compared to using only OakInk-Shape data for training, adding GRAB data and object shapes from Object-only datasets both improves the generalization qualities. The improvement from extra GRAB data is easy to explain as training with more HOI data improves the generalizability of the model while it is interesting that adding the object-only datasets also boosts the performance. We believe this is because the extra object data improves the robustness and generalizability of the object encoding. It extends the latent representation expressiveness that the diffusion model learns to generate.
@@
Ablation of unconditional generation quality. The ability to generate both hand and object to form a grasp is another main contribution. We are interested in whether it can benefit from the additional training data as well. We present the ablation study in Table˜3. Similar to the conditional generation ablation study results, the unconditional generation quality also benefits from additional training data consistently. The metrics of Pene. Dep., Intsec. Vol. and Sim. Disp. Mean indicate that the generated hand-object pair improves plausibility along with adding more training data. The increasing user scoring result suggests that the generated object and hand also become more and more visually realistic along with adding more training data.
@@
Ablation of synchronous denoising schedulers. As one of the main implementation innovations of JHOD, we conduct an ablation about Asynchronous Denoising Schedulers in Table˜4. For the fairness of the comparison, we keep the curriculum of the training the same by mixing data samples from different resources. Intuitively, we designed this module to decouple the noise level in hand latent and object latent so that the model could better use the data with heterogeneous annotations and learn to denoise a certain modality. Without Asynchronous Denoising Schedulers, the noise diffusion and denoising time step of the object part is the same as the hand latent code. It makes the training biased to the HOI dataset where both modalities are available for supervision. Adding Asynchronous Denoising Schedulers, the denoising is learned with more independence for the object part and the hand part. The model can better learn object shape coding to generate hand grasps. Therefore, when evaluating on the out-of-domain objects generated by JHOD, the generation quality is improved as expected.
@@ 5 Conclusion
In this work, we propose a joint diffusion model for generating hand-object grasps. By encoding the modalities of objects and hands into a unified latent space, our model can generate grasps both jointly and conditionally. We address the limitations of existing methods that rely on limited hand-object interaction (HOI) data for training and explore strategies for learning grasp generation that generalizes to a wider variety of objects. Our approach leverages data with partial annotations, thereby mitigating the dependency on fully annotated datasets. The generated grasps, in both unconditional and conditional settings, demonstrate high quality. We believe that our method offers a significant advancement in learning generalizable and robust hand-object grasp generation, even with limited fully annotated data.
Appendix A Grasp Diversity
Given a single object shape, we could generate a set of different grasps onto it. Though diversity is expected to be constrained by the hand grasp data available in training, there are some other implicit but more fundamental cues to help generate diverse hand grasp. This underlying but fundamental constraint can be leveraged to enhance the generation robustness over unseen objects if the feature of the object shape is sufficiently generalizable. For example, the model can learn to avoid penetration between the object volume and the hand shape and certain fingers should be close to the surface of objects to form a physically valid grasp. We generate grasps with unseen objects from the OakInk-shape test set as the condition in Figure˜6. We observe some grasp patterns not provided in the training set, for example, holding a dagger between two fingers. As the implicit concept of forming a valid grasp is always conditioned to object shapes, our strategy of exposing the model training to more diverse object shapes can help to learn grasp patterns on unseen and even out-of-domain objects. We believe this is a key reason for making our method outstanding when generating grasps on out-of-domain objects.
@@ Appendix A Grasp Diversity
Appendix B Dataset Statistics
Here we provide more details about the statistics of the datasets involved in our training and evaluation in Table˜5. The HOI datasets with deformable and articulable hand models, i.e., MANO (Romero et al., 2022), face severe limitations of object resources. Combining the training set of GRAB (Taheri et al., 2020) and OakInk-Image (Yang et al., 2022) datasets still make just $\sim$100 objects. On the other hand, AffordPose (Jian et al., 2023) and DexGraspNet (Wang et al., 2023) contain more object shapes but the different choices of hand models make them hard to integrate with MANO-parameterized datasets. Fortunately, we have the large-scale 3D object shape dataset ShapeNet (Chang et al., 2015) with more than 50,000 objects. We combine the objects from these datasets in the training for the object part. They help to construct a more universal prior distribution for object generation and the grasp posterior distribution.
Table 5: Statistics of the datasets used for training. ShapeNet contains the most object assets. However it is a generic object shape dataset, thus many included objects are not proper for hand grasp. OakInk-Shape contains rich grasp and object data but can not provide annotation for object transformation. We use the grasp data from GRAB and the object data from AffordPose as the supplement during training.
Datasets #obj #grasp real/syn. hand model
ShapeNet 51,300 - real -
GRAB (Taheri et al., 2020) 51 1.3k real MANO (Romero et al., 2022)
OakInk-Image  (Yang et al., 2022) 100 49k real MANO (Romero et al., 2022)
OakInk-Shape  (Yang et al., 2022) 1,700 - real + synthetic MANO (Romero et al., 2022)
AffordPose (Jian et al., 2023) 641 26k synthetic GraspIt (Miller and Allen, 2004)
DexGraspNet (Wang et al., 2023) 5,355 1.32M synthetic ShadowHand (6)
Appendix C Grasp-conditioned Image Generation
We demonstrate using JHOD as a creativity tool to generate 3D grasps and the corresponding images. In Figure˜7, we first use JHOD to sample interaction pairs of 3D hands and objects. With the rendering of the 3D hand-object grasp as the input, we applied Adobe Firefly (Adobe, 2024), an existing tool to generate the images with the depth and edge reference of the 3D pairs. The hand pose and the geometry provide guidance for Firefly to complement the details given simple text prompts. In Figure˜8, compared to grasping images authorized purely by text prompts, images guided by JHOD’s output have more consistent gestures, better-aligned object geometry and boundaries, and fewer hand artifacts. Therefore, with the rise of text-to-image generation, our hand-object grasp generation model can serve as a proxy to enhance the consistency among multiple instances and the visual plausibilities. We also provide a closer look at their detailed in Figure˜9 to compare the results with and without the generated grasp as a proxy. At each row, the images are generated from the same text prompt as shown at the bottom. There are two main benefits of generating images conditioned to the grasp. It first allows a set of images with aligned hand and object geometry and boundaries which can be useful for image and video editing. On the other hand, even though the Adobe Firefly generator has shown a significant advantage over the public Stable Diffusion in eliminating artifacts, we still observed many finger artifacts when generating the images without a grasp condition. Fortunately, with the generated grasp, including the object shape and the hand, as the condition, the image generator can produce significantly fewer artifacts, especially artifact fingers, which have been a notorious issue in image generation recently. We provided some zoomed-in examples at the bottom. Moreover, compared to previous works using hand shapes to guide image generation (Ye et al., 2023b; Narasimhaswamy et al., 2024), we could generate both the object and the hand. Therefore, we could use the whole scene of hand-object interaction as the condition and thus enable more controllable details in the generated image.
@@ Appendix C Grasp-conditioned Image Generation
Appendix D Failure cases
@@
To provide a transparent evaluation of our proposed method, we still notice some failure cases on the unseen objects from OakInk and GRAB datasets as shown in Figure˜10. There are in general three patterns of failures: (1) the hand and the object have no contact, making the grasp physically implausible; (2) the hand skin is twisted or the pose is not able to grasp the object in human common sense; (3) the mesh of object and hand has penetration and intersection. There can be some reasons that make these happen. First of all, we do not explicitly supervise the contact between the object surface and hand as this is a rare annotation on many data sources, and calculating it can be computationally expensive. On the other hand, some implausible grasps have pretty good visual quality as shown in the middle two columns in the figure. However, according to our life experience, we know that the grasp is not physically plausible to manipulate the object. This has gone beyond our scope in this work as we do not have any physics-aware supervision such as the physical demonstration in a physics simulator. Finally, we use the object point cloud to represent object shapes to allow grasp generation on universal objects without a template, this causes the potential penetration between object mesh and hand mesh as the object mesh is ambiguous and unknown to our method. Despite the failure cases, we note that they make only a very small portion of the generated samples (less than 10% by a rough estimation). We show them here for transparency and to help discussion about future works in this area.
@@
Appendix E Hand Contact with Object
Hand contact with the object is a key component of a valid hand-object grasp. In Figure˜11, we visualize the contact area on the hand mesh when the hand is posed to grasp an object. We provide the visualization of the hand-object grasp in three different camera views to showcase the full configuration of the grasp, and also a hand-only visualization to more clearly justify the contact area (in red). Following existing practice while making a stricter standard to justify a valid “contact”, we set the distance threshold between the hand mesh face and the object mesh face to be 0.005 meter for the visualization.
Appendix F Qualitative Comparison
In addition to the quantitative results presented in the experiment sections, we now provide qualitative comparisons between our method and the baseline GrabNet (Taheri et al., 2020), as shown in Figure˜12. We randomly sample objects from the "bottle" category in OakInk that were not seen during training. Grasps are generated for each object without any cherry-picking. In each result pair, the left-hand grasp is produced by our method, while the right-hand grasp is from GrabNet. As shown, GrabNet generates grasps that penetrate the object in pairs (1), (2), (4), and (6), and lacks plausible contact in pairs (3) and (8). In contrast, our method produces more realistic grasps in most cases, although some penetration is still observed in pairs (2) and (6). Overall, these comparisons suggest that our method achieves superior qualitative performance in grasp generation.
Figure 11: Visualization of the generated hand-object grasp and the corresponding hand contact area. Here, we set the distance threshold to justify the contact between a hand mesh face and an object mesh face to be 0.005 meters. The contact area is rendered in red.
Figure 12: Grasps generated by our proposed method and GrabNet (Taheri et al., 2020). In each pair of grasps, the left-hand grasp is generated by our method and the right-hand one is generated by GrabNet.